# Towards a Multi-Embodied Grasping Agent

### v1

@@ VI-C Multi-Embodiment Grasping
Our evaluation protocol is similar to that of [1]. For each gripper, we measure the grasp success rate across 10 held-out test scenes, executing 100 grasps sampled from our model in each scene. As a pre-processing step, we filter out any grasps that are in collision with the scene. However, grasps that result in gripper self-collisions are counted as failures, as such configurations would damage the physical hardware. Our optimized implementation can sample a full batch of 100 grasps for a dexterous gripper in under 10 seconds on a single consumer-grade GPU.
@@
For the single-embodiment benchmark, we train a dedicated model for each gripper type on a single H200 GPU for 500 scene-epochs. We use a batch size of five scenes, from which we sample 128 grasps per batch. Note that since scenes can contain up to 1,500 valid grasps, each epoch consists of a different random subset. We employ several data augmentation techniques: randomizing the Farthest Point Sampling (FPS) initialization, applying random rotations around the z-axis, and weighting the loss to emphasize later time steps in the flow, which are critical for refining the final pose. As noted in [6], such rotational augmentations are beneficial even for equivariant models to mitigate the accumulation of minor numerical errors across layers.
@@
For the multi-embodiment task, we train a single, unified model using a total batch size of 10 scenes distributed across all five gripper types, training for the same number of epochs. To improve conditioning, we incorporate a technique from Lim et al [37] by adding a sixth ‘dummy’ gripper with zero DoFs. This dummy gripper serves as an unconditional class, which helps steer the flow generation process more effectively for the specific, conditioned gripper types. During training, a fraction of the data from all real grippers is repurposed for this dummy class by replacing the true gripper label and nullifying the joint information. The results, presented in Table III, indicate a clear performance advantage for the generalist multi-embodiment model in four out of five gripper types.
@@ VII Limitations and Future Work
A further limitation stems from an architectural trade-off made to simplify the model. In contrast to MultiDiffusion [1], our current approach does not include an explicit geometric representation of the gripper’s body during grasp synthesis. This simplification leads to an increased rate of collisions between the gripper and the scene. While we currently filter these invalid grasps in a post-processing step, efficiently re-integrating the full gripper geometry into the model is a key priority for future work.
A prerequisite for our method is a JAX [10] implementation of the gripper’s kinematic model. While we implemented these manually, this process could be largely automated by developing a script to parse standard URDF files, facilitating the integration of new gripper types. Our geometric joint representation can naturally accommodate under-actuation via masking during the feature readout; however, we leave the implementation of this capability for future work. As a result, joint constraints are currently enforced softly, learned implicitly from the training data. Experiments with partial point cloud observations were omitted due to increased engineering efforts with variable sized point clouds in JAX [10].
We optimized each module using JAX’s vector mapping capabilities to enable parallelization across scenes, grippers, and grasps. As a result, our method can generate a batch of 100 grasps for a dexterous gripper in under 10 seconds on a consumer-grade GPU (<<12GB VRAM), a task previously infeasible with comparable Torch-based implementations [5]. Although equivariant methods have a reputation for being memory-intensive, some of our optimized modules are already highly efficient. For instance, the Kinematic Encoder processes a batch of 128 configurations over all five grippers i.e. 640 grasps configurations total, in under 10 ms for one training iteration. This level of performance suggests that such geometric encoders are becoming viable for latency-sensitive tasks beyond grasping. While our full model is not yet ready for real-time online applications, future work could leverage low-level equivariant operations implemented in CUDA e.g. cuEquivariance to bridge this gap. Such efforts are well-justified, as the promise of these methods extends beyond mere efficiency; a strong argument for their use in manipulation is their native capacity to represent vector-based information, such as friction forces and surface normals, while seamlessly integrating scalar features from color or foundation models.

### v2

@@ VI-C Multi-Embodiment Grasping
For each gripper, we measure the grasp success rate and joint diversity across 10 held-out test scenes, executing 100 grasps sampled from our model in each scene. During pre-processing, we filter out any grasps that are in collision with the scene. Our optimized implementation can sample a full batch of 100 grasps for a dexterous gripper in under 10 seconds on a single consumer-grade GPU.
@@
For the single-embodiment benchmark, we train a dedicated model for each gripper type on a single H200 GPU for up to 500 scene-epochs. We use a batch size of five scenes, from which we sample 128 grasps per batch. Note that since scenes can contain up to 1,500 valid grasps, each epoch consists of a different random subset. We employ several data augmentation techniques: randomizing the Farthest Point Sampling (FPS) initialization, applying random rotations around the z-axis, and over-sample later time steps in the flow, which are critical for refining the final pose. As noted in [5], rotational augmentations are beneficial even for equivariant models to mitigate the accumulation of minor numerical errors across layers.
@@
For the multi-embodiment task, we train a single, unified model with a same-sized scene batch count distributed across all five gripper types. To improve conditioning, we incorporate a technique from Lim et al [37] by adding a sixth ‘dummy’ gripper with zero DoFs. This dummy gripper serves as an unconditional class, which helps steer the flow generation process more effectively for the specific, conditioned gripper types. During training, a fraction of the data from all grippers is repurposed for this dummy class by replacing the true gripper label and nullifying the joint information. To compute a standard diversity metric that is applicable for all gripper types including parallel jaw, we deviate from the common angle variance metric [26] in favor of the mean normalized joint deviation (NJD). This metric normalizes all joint spaces to between plus and minus one and computes the mean of the standard deviations over all DoFs. The results are presented in Table IV. Especially for the humanoid grippers we observe an increase in both grasp success rate and diversity, indicating an advantageous transfer between datasets. The DEX-EE gripper does not seem to profit in terms of grasp success rate, but upon inspection shows an overall more diverse grasp coverage. The parallel-jaw grippers’ performance remains nearly unchanged, as both single and multi-embodiment models saturate the grasp success metrics, while showing more diverse finger placement for the ViperX gripper.
In a second experiment, we perform a direct comparison against a re-implemented version of Diffusion [10]. To ensure a fair comparison and isolate the impact of the core learning objective, we use the same encoder pipeline for both methods, holding all other architectural and training parameters constant. We generated a new dataset for this experiment by restricting it to the static pre-grasp configurations required by the baseline; for the Shadow Hand, we used the three-fingered configuration from the original work [10]. Since it is only used for this benchmark, we reduced this dataset to only three representative gripper types and generated 15 000 scenes in total. Accordingly, our model was adapted to this setting by providing the fixed joint configurations as static inputs. The results are presented in Table III. While we do not observe a clear advantage of one method over the other in this benchmark, diffusion-based approaches are significantly harder to deploy for the general case where all gripper DoFs are learned through geometric representations. Given a uniform distribution over the full action space, the flow-based approach respects action bounds within physically plausible limits throughout the process. Violating this requirement, as in diffusion-based approaches, results in physically impossible joint settings and rotations exceeding 360 degrees, which the model cannot distinguish from valid gripper settings, resulting in training instabilities.
@@ VII Limitations and Future Work
A further limitation stems from an architectural trade-off made to simplify the model. In contrast to Diffusion [10], our current approach does not include an explicit geometric representation of the gripper’s body during grasp synthesis. This simplification leads to an increased rate of collisions between the gripper and the scene. While we currently filter these invalid grasps in a post-processing step, efficiently re-integrating the full gripper geometry into the model is a key priority for future work. Our geometric joint representation can naturally accommodate under-actuation via masking during the feature readout; however, we leave the implementation of this capability for future work. As a result, joint constraints are currently enforced softly, learned implicitly from the training data. Experiments with partial point cloud observations were omitted due to increased engineering efforts with variable sized point clouds in JAX [9]. We optimized each module using JAX’s vector-mapping capabilities to enable parallelization across scenes, grippers, and grasps. As a result, our method can generate a batch of 100 grasps for a dexterous gripper in under 10 seconds on a consumer-grade GPU (<<12GB VRAM), a task previously infeasible with comparable Torch-based implementations [4]. Although equivariant methods have a reputation for being memory-intensive, some of our optimized modules are already highly efficient. For instance, the Kinematic Encoder processes a batch of 128 configurations over all five grippers i.e., 640 grasps configurations total, in under 10 ms for one training iteration. This level of performance suggests that such geometric encoders are becoming viable for latency-sensitive tasks beyond grasping. Such efforts are well-justified, as the promise of equivariant methods extends beyond mere efficiency; a strong argument for their use in manipulation is their native capacity to represent vector-based information, such as friction forces and surface normals, while seamlessly integrating scalar features from color or foundation models.